name the physical quantity which is expressed as the dot product of force and velocity  its a scalar or vector quantity

Grade:11

3 Answers

Khimraj
6 years ago
The physical quantity obtained by dot product of force and velocity is Power. It is Scalar quantity.
F.v = F.d/t = W/t = P.
Hope it clears. If you have doubt you can ask. If you like anser then please approve it. 
Arun
6 years ago
If It is Force times Velcoity (F*V) then that equals Power. Power = Fv. Thus Power can be expressed in terms of Force times Velocity. And Also it is a Scalar Quantity. Because a vector times a vector is = a scalar. P.S :-- Force * Velocity = Force * Displacement / Time (Force*Displacement = Work) = Work/Time = Power.
fatima
5 years ago
electric flux which is a dot product of electric feild intensity and area of flux these both are vectors but we are getting scaler the electric flux